OpenStudy (anonymous):

Input in standard form the equation of the given line. The line that passes through (1, 5) and (-2, 3)

OpenStudy (anonymous):

Standard form is Ax + By = C

hero (hero):

y = mx + b 5 = 1m + b 3 = -2m + b 5 - m = b 3 + 2m = b b = b 5 - m = 3 + 2m 5 - 3 = 2m + m 2 = 3m 2/3 = m b = 3 + 2m b = 3 + 2(2/3) b = 3 + 4/3 b = 9/3 + 4/3 b = 13/3 y = (2/3)x + 13/3 y = (2x + 13)/3 3y = 2x + 13 0 = 2x - 3y + 13 -13 = 2x - 3y Thus the standard form of the line is: 2x - 3y = -13

OpenStudy (anonymous):

Great work!

hero (hero):

Okay, you should be able to figure them out on your own from here.

hero (hero):

@Sephichief